OpenStudy (anonymous):

x = y2 - 1 2x - y = 4 (3, ) and (, )

OpenStudy (anonymous):

quadratic system of equations

mathslover (mathslover):

i didn't understand the question completely wait

OpenStudy (anonymous):

i can post the attachment

mathslover (mathslover):

please

OpenStudy (anonymous):

^^

mathslover (mathslover):

So you are giving test right ?

OpenStudy (anonymous):

no this is an assignment

mathslover (mathslover):

hmn what you tried ?

OpenStudy (anonymous):

well I tried substitution, eliminations,quadratic formula, foiling...anything else and I would lose my mind

OpenStudy (anonymous):

foiling sounds like fun...

OpenStudy (anonymous):

lol yea FUN!!

OpenStudy (anonymous):

xD

mathslover (mathslover):

ok so u r having problem in steps ... let me give u 1st stem .. : \(\huge{x=y^2-1}\) let this equation be 1st .. and similarly 2x-y be the 2nd

mathslover (mathslover):

put the value of x in 2nd one.. what u get ?

OpenStudy (anonymous):

2(y^2-1)-y=4

mathslover (mathslover):

Gr8 now simplify it by opening brackets

OpenStudy (anonymous):

2y^2-1-y=4

mathslover (mathslover):

right can u write it as \(2y^2-y-1=4\) ?

OpenStudy (anonymous):

ok \[2y^2-y-1=4\]

mathslover (mathslover):

right transpose -1 to Right Hand Side ( RHS )

OpenStudy (anonymous):

\[2y^2-y=5\]

mathslover (mathslover):

right : now transpose 5 to Left Hand Side again :

OpenStudy (anonymous):

\[2y^2-y-5\]

mathslover (mathslover):

= ?

OpenStudy (anonymous):

\[2y^2-y-5=0\]

mathslover (mathslover):

right. .. what type of equation is this ?

OpenStudy (anonymous):

quadratic

mathslover (mathslover):

solve this by the following formula : \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

mathslover (mathslover):

a = 2 , b= -1 , c=-5

OpenStudy (anonymous):

i know im working it out srry if im taking to long

mathslover (mathslover):

no problem @aussy123 u can take ur time as much u want

OpenStudy (anonymous):

\[-1\pm\frac{\sqrt{-41}}{4}\\]

OpenStudy (anonymous):

\[-1\pm\frac{\sqrt{-41}}{4}\\] \]

mathslover (mathslover):

that's right

mathslover (mathslover):

wait but what does (3, ) mena

OpenStudy (anonymous):

thats the coordinates they are looking for

mathslover (mathslover):

wait i am calling some body good at this... or i am trying...

OpenStudy (anonymous):

okay

mathslover (mathslover):

@dpaInc

mathslover (mathslover):

(3,2) and (5/4 , -3/2) is what i got

OpenStudy (anonymous):

how

mathslover (mathslover):

wait

mathslover (mathslover):

u did wrong in some steps ... \[2(y^2-1)-y=4\] \[2y^2-2-y=4\] \[2y^2-y=6\] \[2y^2-y-6=0\] \[\textbf{a=2 , b = -1 , c= -6}\]

OpenStudy (anonymous):

oh ok I see lol

mathslover (mathslover):

so shall i think that u have got the answer ?

OpenStudy (anonymous):

yea